OpenStudy (anonymous):

a3+b3 varify by an activity

jhonyy9 (jhonyy9):

a^3 +b^3 =(a+b)(a^2 -ab +b^2) =a^3 -a^2b +ab^2 +a^2b -ab^2 +b^3 =a^3 +b^3 you have thought it hence ? good luck bye

OpenStudy (anonymous):

this is not true ans. of my question

OpenStudy (accessdenied):

Can you explain the question? I don't even understand what we're trying to 'verify' here. The formula, a^3 + b^3 = (a + b)(a^2 - ab + b^2)? Are we supposed to make some 'activity' that proves that the formula is true?

OpenStudy (anonymous):

a]find the area of any triangle, b]find the area of any polygon by completing the rectangle. c]obtain a square on a given line segment, d]given an area,obtain different polygons of the same area. 15}to verify that the angle in the same segment of a circle are equal,using the method of paper cutting,pasting,and folding. 18}to verify the identity a3-b3=(a-b)(a2 ab b2) ,for simple cases using a set of unit cubes. 19}to verify the identity a3 b3=(a b)(a2-ab b2) ,for simple cases using a set of unit cubes. 20}to verify the identity (a b)3=a3 b3 3ab(a b) ,for simple cases using a set of unit cubes. 21}to verify the identity (a-b)3=a3-b3-3ab(a-b) ,for simple cases using a set of unit cubes. 22}to interpret geometrically the factor of a quadratic expression of the type x2 bx c,using square grids,strips and paper slips.”

OpenStudy (anonymous):

refer to this link and check out task 2 in it ........ it will surely help you]